# FreeRide: Harvesting Bubbles in Pipeline Parallelism

### v1

@@ Abstract.
The occurrence of bubbles in pipeline parallelism is an inherent limitation that can account for more than 40% of the large language model (LLM) training time and is one of the main reasons for the underutilization of GPU resources in LLM training. Harvesting these bubbles for GPU side tasks can increase resource utilization and reduce training costs but comes with challenges. First, because bubbles are discontinuous with various shapes, programming side tasks becomes difficult while requiring excessive engineering effort. Second, a side task can compete with pipeline training for GPU resources and incur significant overhead. To address these challenges, we propose FreeRide, a system designed to harvest bubbles in pipeline parallelism for side tasks. FreeRide provides programmers with interfaces to implement side tasks easily, manages bubbles and side tasks during pipeline training, and controls access to GPU resources by side tasks to reduce overhead. We demonstrate that FreeRide achieves 7.8% average cost savings with a negligible overhead of about 1% in training LLMs while serving model training, graph analytics, and image processing side tasks.
@@ 1. Introduction
In this paper, we present FreeRide, a holistic system to harvest bubbles in pipeline parallelism to serve extra GPU workloads as side tasks. There are two main challenges that FreeRide has to overcome. First, customizing side tasks for bubbles of various shapes, i.e., their duration and available GPU memory, requires enormous programming effort. Second, LLM training requires high-end GPUs that are expensive and in high demand. If side tasks interfere with the main pipeline training workload, e.g., accessing more GPU resources than bubbles can provide, they will slow down pipeline training and significantly increase training costs.
@@
Our approach to overcoming the programming complexity is based on the observation that many GPU workloads naturally consist of small, repetitive steps, such as the epochs in model training that repeatedly load data and update model weights. To reduce the programming effort, FreeRide introduces a framework that abstracts away the implementation details of side tasks, allowing programmers to adapt various side tasks to fit into the bubbles. The key idea is to represent the life cycle of a side task, from its process creation to termination, as states in a state machine. FreeRide provides two sets of unified interfaces — the iterative interface that features lower performance overhead, and the imperative interface that features better versatility. They facilitate the implementation of side tasks as state transitions with little engineering effort. FreeRide manages side tasks through these interfaces and serves them during bubbles.
@@
FreeRide limits the GPU resource consumption of side tasks through the automated side task profiler and the side task manager. The side task profiler first captures the key performance characteristics of the newly implemented side tasks. The side task manager coordinates a group of side task workers, one for each GPU in the platform, and assigns each of the side tasks to one of the workers based on the characteristics. During pipeline training, bubbles are reported to the side task manager from the FreeRide-instrumented pipeline training system. The side task manager starts side tasks when the bubble period begins and pauses them when the bubble ends. A side task worker deploys each side task on top of CUDA MPS (nvidia_developers_multi-process_2024, ) to limit its GPU memory consumption and uses a containerized environment, e.g., Docker (bernstein_containers_2014, ) for isolation. These components work collaboratively to serve side tasks during bubbles, achieving a low performance overhead on the primary pipeline training workload.
@@
In summary, FreeRide is a holistic solution that manages and serves the side task by leveraging bubbles in pipeline training, while maintaining minimal performance overhead and requiring low programming effort. We evaluate FreeRide by deploying it to run side tasks alongside DeepSpeed that runs pipeline training (rasley_deepspeed_2020, ). We measure the time increase of pipeline training as the performance overhead caused by side tasks. As the throughput of different side tasks is not directly comparable with the pipeline training workload, we use the cost of GPUs as a unified metric, i.e., the cost of the extra execution time from co-locating side tasks with pipeline training vs. the cost saved from running side tasks that otherwise would run on dedicated lower-tier GPUs.
@@
  • We present the programming framework and interfaces of FreeRide based on a state machine abstraction to implement generic side tasks with little engineering effort.
@@
  • By serving side tasks based on the iterative interface, FreeRide achieves average cost savings of 7.8% with a low performance overhead of 1.1%. This is significantly better than using CUDA MPS (nvidia_developers_multi-process_2024, ) directly to co-locate the tasks, which results in a 4.5% cost increase and 48.7% overhead. When handling a mix of these 3 types of side tasks, FreeRide achieves 10.1% cost savings with a 1.1% overhead.
@@ 2.1. Pipeline Parallelism and Bubbles
There are periods in pipeline training when the estimated GPU streaming multiprocessor (SM) occupancy is low, as depicted by the green curves in Figure 1(a). We refer to these periods as bubbles in the pipeline, marked as shaded areas. Bubbles inherently exist in pipeline parallelism and occur repetitively throughout training, as they are fundamentally caused by unsatisfied dependencies between FP and BP operations (li_chimera_2021, ; liu_hanayo_2023, ). In the example of Figure 1, Stage 1 must wait for input from Stage 0 before it can execute its first FP operation, creating a bubble in Stage 1 that starts from $t+0$.
@@ 2.2.1. Bubble Categorization
We categorize the bubbles into 3 types based on their positions in the training pipeline and their causes.
@@
We further study pipeline training of models of different sizes. As shown in Figure 2(a), bubble shapes differ. Overall, bubbles in larger LLMs have less available memory and shorter duration, but the distributions are similar as training follows the same pipeline schedule. Even larger models do not eliminate bubbles as they inherently exist. Under the same configuration, the characteristics of bubbles remain the same during training as epochs are repetitive and stable.
Figure 1. A pipeline training epoch in DeepSpeed.
[…]
@@ 2.3. Utilizing Bubbles
The difficulties in mitigating bubbles in pipeline parallelism motivate an alternative approach — acknowledging their existence and leveraging their resources by allocating additional GPU workloads to them. Prior work has utilized bubbles to run procedures that enhance pipeline training. For example, Bamboo uses bubbles to perform redundant computation for successive layers to improve the reliability of pipeline training on spot instances (thorpe_bamboo_2023, ); PipeFisher computes second-order optimization based on the Fisher information matrix to speed up convergence (osawa_pipefisher_2023, ). However, they tightly couple the pipeline training system with the specialized procedures. Implementing the specialized procedures is complicated, especially since such customization should consider various bubble shapes — with durations ranging from 0.22 to 1.04 seconds, and available GPU memory from less than 3 GB to more than 20 GB on each GPU (Section 2.2).
@@ 2.4. Challenges
Challenge 1: programming effort required to implement side tasks. Typically, GPU workloads are implemented based on the assumption that they have access to the full GPU and can run continuously until they finish execution. However, bubbles are intermittent and largely vary in duration, as discussed in Section 2.2. A side task should be tailored to bubble patterns — the side task automatically pauses or resumes when a bubble ends or starts. Customizing the training framework to embed side tasks is conceptually feasible but limits the flexibility of implementing and executing generic GPU workloads, much like the limitations from prior work on co-running specialized procedures (osawa_pipefisher_2023, ; thorpe_bamboo_2023, ).
@@ 3. FreeRide Design Overview
FreeRide is our system that addresses the aforementioned challenges in utilizing bubbles in pipeline training to serve generic GPU side tasks. FreeRide minimizes the performance impact of side tasks on pipeline training. In this section, we present the design and high-level ideas of FreeRide.
@@ 3.1. Side Task Programming Interface
Given the high cost and priority of the main pipeline training workload, the side task should not overlap with this main task so as to avoid competing for GPU resources. This requirement is challenging from a programmer’s perspective as it is difficult to tailor every workload to different bubble shapes. We observe that GPU workloads are not monolithic, rather, they can be often divided into small, repeated steps with largely predictable per-step duration. For example, epochs in model training, iterations in graph analytical workloads (xu_gardenia_2019, ; page_pagerank_1999, ; koren_matrix_2009, ), and steps to process each image in image-processing workloads (nvidia_developers_image_2019, ) all follow this pattern. On the other hand, bubbles also demonstrate repeating and predictable patterns, as discussed in Section 2.2.
With these observations in mind, our idea is to provide an iterative programming interface that supports the step-by-step execution of side tasks to fit bubble patterns — the side tasks can be naturally divided into smaller steps to fit into bubbles of different shapes. This interface employs a state machine abstraction for the life cycle of a side task composed of different states during its execution. The execution of side tasks within bubbles can be implemented as state transition functions in Figure 4 (details in Section 4.1). An execution pattern tailored to control the start and pause of side tasks through state transitions minimizes the performance impact on pipeline training.
@@ 3.2. Profiling-guided Side Task Management
To judiciously manage side tasks on bubbles, FreeRide leverages offline profiling to understand the shapes of bubbles. Then, when a side task is newly submitted to FreeRide, as shown in Figure 3, FreeRide’s automated side task profiler tracks its GPU memory consumption and per-step duration. During execution time, FreeRide employs one side task manager and multiple side task workers, one for each GPU. The side task manager assigns the newly submitted side task to one of the side task workers which will create the side task process, based on the resulting profile. We instrument DeepSpeed to report the start timestamps and duration of bubbles to the side task manager that will initiate state transitions of each side task through remote procedure calls (RPCs) at the start and end of each bubble.
@@ 3.3. FreeRide Workflow
Putting the aforementioned ideas together, we present the workflow of FreeRide in Figure 3. First, programmers adapt their side task implementation using the programming framework provided by FreeRide (step ➊). FreeRide then automatically generates a profile of the side task’s characteristics (step ➋), which is submitted with the side task to the side task manager of FreeRide (step ➌). After the side task is submitted for execution, based on the memory allocation of pipeline training and the characteristics of the side task, the side task manager will assign this side task to one of the workers (step ➍). When the main pipeline training workload is running, the side task manager continuously adds bubbles from the instrumented training framework (step ➎); at the same time, it starts/pauses side tasks based on the available bubbles (step ➏).
@@ 4.1. Programming Framework of FreeRide
Figure 4. State transitions in a side task program.
[…]
Figure 4(a) describes the programming framework of a side task. The framework’s core is a state machine with five states and six state transitions. These five states capture the life cycle of a side task, from process creation to process termination, and correspond to different usage of hardware resources, e.g., GPU memory and GPU execution time. The six state transitions are used by the programmer to implement the user-defined logic of a side task. The programmer can override the state transition functions to define what happens during state transitions, e.g., allocating or releasing hardware resources or performing computation on GPU. Once the side task is implemented, FreeRide automatically handles the state transitions at runtime. Next, we discuss the states and their transitions.
@@
$\bullet$ RUNNING. In this state, the side task executes the step-wise GPU workload. Referring to the example above of the model training side task, this step involves reading the next batch, computing the output and loss, updating the model weights, and resetting the optimizer states. The side task iteratively enters the RunNextStep() state transition to execute these steps until the side task manager transitions its state through PauseSideTask(). Therefore, in this state, the side task process uses both the GPU memory and the GPU SMs.
$\bullet$ STOPPED. This state marks the end of the life cycle of a side task, where the side task process releases all of its hardware resources and terminates. It can be transited from states CREATED, PAUSED, and RUNNING through StopSideTask() initiated by the side task manager.
Figure 4(b) shows state transitions of a side task in Stage 0 of Figure 1. Initially, the side task is in the PAUSED (P) state. After four FP operations in the main training workload have finished, a bubble starts and the side task manager initiates StartSideTask() to transit the side task to the RUNNING (R) state. After the first bubble ends, the side task manager pauses the side task via PauseSideTask(). Then, the main training workload has BP operations and bubbles interleaved, leading to back-and-forth transitions between PAUSED and RUNNING states of the side task.
@@ 4.2. Interface for Side Task Implementation
Iterative programming interface. This is the preferred interface for side tasks in FreeRide. It periodically checks whether the side task manager has initiated any state transitions. If so, it executes the state transition functions in Figure 4(a) and updates the state of the side task. Then, if the side task is currently in the RUNNING state, it executes RunNextStep(). The programmer only has to override these transition functions to implement the side task. Pausing and resuming the side task, the transition of states, and communication with the FreeRide side task manager are all handled by the interface itself. GPU workloads that are naturally step-wise, e.g., model training, can be easily adapted to the iterative interface. We will discuss the adaption to the iterative interface in Section 5 using an example.
Imperative programming interface. Not all side tasks can be explicitly implemented step-wise. Therefore, FreeRide provides the imperative interface as a fallback solution. The core is the function RunGpuWorkload() that allows the programmer to implement generic GPU side tasks without breaking them into steps. When the side task manager transits the state of the side task to RUNNING for the first time, the interface calls the RunGpuWorkload() function to execute the side task. The interface implements the pausing and resuming through signals (SIGTSTP and SIGCONT (job_control_signals, )) and calls StartSideTask() and PauseSideTask() inside the handlers of the two signals. The imperative interface offers better versatility at the cost of higher performance overhead (discussed in Section 5 and evaluated in Section 6.2).

### v2

@@ Abstract.
The occurrence of bubbles in pipeline parallelism is an inherent limitation that can account for more than 40% of the large language model (LLM) long training times and is one of the main reasons for the under-utilization of GPU resources in LLM training. Harvesting these bubbles for GPU side tasks can increase resource utilization and reduce training costs but comes with challenges. First, because bubbles are discontinuous with various shapes, programming side tasks becomes difficult while requiring excessive engineering effort. Second, a side task can compete with pipeline training for GPU resources and incur significant overhead. To address these challenges, we propose FreeRide, a middleware system that harvests the hard-to-utilize bubbles in pipeline parallelism systems to run generic GPU side tasks. FreeRide provides programmers with interfaces to implement side tasks easily, manages bubbles and side tasks during pipeline training, and controls access to GPU resources by side tasks to reduce overhead. We demonstrate that FreeRide achieves almost 8% average cost savings with a negligible overhead of about 1% in training LLMs while serving model training, graph analytics, and image processing side tasks.
Large language model, pipeline parallelism, utilization, GPU system
journalyear: 2025copyright: acmlicensedconference: 26th ACM Middleware Conference; December 15–19, 2025; Nashville, TN, USAbooktitle: 26th ACM Middleware Conference (Middleware ’25), December 15–19, 2025, Nashville, TN, USAdoi: 10.1145/3721462.3730950isbn: 979-8-4007-1554-9/25/12submissionid: 16ccs: Computing methodologies Machine learningccs: Computer systems organization
@@ 1. Introduction
In this paper, we present FreeRide, a middleware system that bridges the gap between the available yet hard-to-utilize bubbles in pipeline parallelism and the extra GPU workloads we run to harvest them. We refer to these extra GPU workloads as side tasks. There are two main challenges that FreeRide has to overcome. The first challenge is the programming complexity. Bubbles are of various shapes, i.e., their duration and available GPU memory. Customizing side tasks for these bubbles by doing ad-hoc implementation requires enormous programming effort. Second, LLM training requires high-end GPUs that are expensive and in high demand. If side tasks interfere with the main pipeline training workload, e.g., overlapping their GPU execution with pipeline training or accessing more GPU resources than bubbles can provide, they will slow down pipeline training and increase training costs.
@@
Our approach to overcoming the programming complexity is based on the observation that many GPU workloads naturally consist of small, repetitive steps, such as the epochs in model training that repeatedly load data and update model weights. FreeRide operates between the pipeline parallel training and the generic GPU side tasks implemented by the user. To reduce the programming effort, FreeRide introduces a framework that abstracts away the implementation details of side tasks, allowing programmers to adapt various side tasks to fit into the bubbles. The key idea is to represent the life cycle of a side task, from its process creation to termination, as states in a state machine. FreeRide provides two sets of unified interfaces — the iterative interface that features lower performance overhead, and the imperative interface that features better versatility. They facilitate the implementation of side tasks as state transitions with little engineering effort. FreeRide manages side tasks through these interfaces and serves them during bubbles.
@@
FreeRide limits the GPU resource consumption of side tasks through the automated side task profiler and the side task manager. The side task profiler first captures the key performance characteristics of the newly implemented side tasks. The side task manager coordinates a group of side task workers, one for each GPU in the platform, and assigns each of the side tasks to one of the workers based on the characteristics. During pipeline training, bubbles are reported to the side task manager from the FreeRide-instrumented pipeline training system. The side task manager starts side tasks when the bubble period begins and pauses them when the bubble ends. A side task worker deploys each side task on top of CUDA MPS, which enables the concurrent execution of CUDA kernels from different processes (nvidia_developers_multi-process_2024, ) to limit the side task’s GPU memory consumption and uses a containerized environment, e.g., Docker (bernstein_containers_2014, ) for isolation. These components work collaboratively to serve side tasks during bubbles, achieving a low performance overhead on the primary pipeline training workload.
@@
In summary, FreeRide is a middleware system that bridges the gap between the resourceful yet hard-to-utilize bubbles in pipeline parallelism and the extra GPU workloads we run to harvest the bubbles. It provides a holistic solution to manage and serve side tasks by leveraging bubbles in pipeline training, while maintaining minimal performance overhead and requiring low programming effort. We evaluate FreeRide by deploying it to run side tasks alongside DeepSpeed that runs pipeline training (rasley_deepspeed_2020, ). We measure the time increase in pipeline training as the performance overhead caused by side tasks. As the throughput of different side tasks is not directly comparable with the pipeline training workload, we use the cost of GPUs as a unified metric, i.e., the cost of the extra execution time from co-locating side tasks with pipeline training vs. the cost saved from running side tasks that otherwise would run on dedicated lower-tier GPUs.
@@
  • We present the programming framework and the interfaces of FreeRide [^1] https://github.com/jiashu-z/freeride based on a state machine abstraction to implement generic side tasks with little engineering effort.
@@
  • By serving side tasks based on the iterative interface, FreeRide achieves average cost savings of 7.8% with a low performance overhead (time increase in pipeline parallel training) of 1.1%. This is significantly better than using CUDA MPS (nvidia_developers_multi-process_2024, ) directly to co-locate the tasks, which results in a 4.5% cost increase and 48.7% overhead. When handling a mix of these three types of side tasks, FreeRide achieves 10.1% cost savings with a 1.1% overhead.
@@ 2.1. Pipeline Parallelism and Bubbles
Figure 1. A pipeline training epoch in DeepSpeed.
[…]
There are periods in pipeline training when the GPU streaming multiprocessor (SM) occupancy is low, as depicted by the green curves in Figure 1(a). We refer to these periods as bubbles in the pipeline, which are marked as shaded areas. Bubbles inherently exist in pipeline parallelism and occur repetitively throughout training, as they are fundamentally caused by unsatisfied dependencies between FP and BP operations (li_chimera_2021, ; liu_hanayo_2023, ). In the example of Figure 1, Stage 1 must wait for input from Stage 0 before it can execute its first FP operation, creating a bubble in Stage 1 that starts from $t+0$.
@@ 2.2.1. Bubble Categorization
We categorize bubbles into 3 types based on their positions in the training pipeline and their causes.
@@
We further study pipeline training of models of different sizes. As shown in Figure 2(a), bubble shapes differ. Overall, bubbles in larger LLMs have less available memory and shorter duration, but the distributions are similar as training follows the same pipeline schedule. Even larger models do not eliminate the inherently exist bubbles. Under the same configuration, bubbles have the same characteristics during training, as epochs are repetitive and stable.
@@ 2.3. Utilizing Bubbles
The difficulties in mitigating bubbles in pipeline parallelism motivate an alternative approach — acknowledging their existence and leveraging their resources by allocating additional GPU workloads to them. Prior work has utilized bubbles to run procedures that enhance pipeline training. For example, Bamboo uses bubbles to perform redundant computation for successive layers to improve the reliability of pipeline training on spot instances (thorpe_bamboo_2023, ); PipeFisher computes second-order optimization based on the Fisher information matrix to speed up convergence (osawa_pipefisher_2023, ). However, they tightly couple the pipeline training system with specialized procedures. Implementing specialized procedures is complicated, especially since such customization should consider various bubble shapes — with durations ranging from 0.22 to 1.04 seconds, and available GPU memory from less than 3 GB to more than 20 GB on each GPU (Section 2.2).
@@ 2.4. Challenges
Challenge 1: programming effort required to implement side tasks. Typically, GPU workloads are implemented assuming that they have access to the full GPU and can run continuously until they finish execution. However, bubbles are intermittent and largely vary in duration, as discussed in Section 2.2. A side task should be tailored to bubble patterns — the side task automatically pauses or resumes when a bubble ends or starts. Customizing the training framework to embed side tasks is conceptually feasible but limits the flexibility of implementing and executing generic GPU workloads, much like the limitations from prior work on co-running specialized procedures (osawa_pipefisher_2023, ; thorpe_bamboo_2023, ).
@@ 3. FreeRide Design Overview
FreeRide is our middleware system that addresses the aforementioned challenges in utilizing bubbles in pipeline training to serve generic GPU side tasks. It minimizes the performance impact of side tasks on pipeline training. In this section, we present the high-level ideas of FreeRide.
@@ 3.1. Side Task Programming Interface
Given the high cost and priority of the main pipeline training workload, the side task should not overlap with this main task to avoid competing for GPU resources. This requirement is challenging from a programmer’s perspective, as it is difficult to tailor every workload to different bubble shapes. We observe that GPU workloads are not monolithic, rather, they can be often divided into small, repeated steps with largely predictable per-step duration and resource consumption, i.e., GPU memory. For example, epochs in model training, iterations in graph analytical workloads (xu_gardenia_2019, ; page_pagerank_1999, ; koren_matrix_2009, ), and steps to process each image in image-processing workloads (nvidia_developers_image_2019, ) all follow this pattern. On the other hand, bubbles also demonstrate repeating and predictable patterns, as discussed in Section 2.2.
With these observations in mind, our idea is to provide an iterative programming interface that can incorporate the step-by-step execution of side tasks to bubbles with various shapes. The user only has to implement the side task step without being concerned with the bubble shapes, and the bubbles can serve these side tasks with largely predictable durations to avoid lack of GPU resource or overlapping of side task execution and pipeline training. The iterative programming interface provided by FreeRide employs a state machine abstraction for the life cycle of a side task composed of different states during its execution. The execution of side tasks within bubbles can be implemented as state transition functions in Figure 4 (details in Section 4.1). FreeRide works as the middleware layer in between the side tasks and the bubbles of pipeline training, managing the side tasks’ start and pause through controlling their state transitions. In this way, FreeRide fits the side tasks into bubbles and minimizes the performance impact on pipeline training.
@@ 3.2. Profiling-guided Side Task Management
To judiciously manage side tasks on bubbles, FreeRide leverages profiling to understand the shapes of bubbles, which can be done offline or during the first few epochs of pipeline training. Then, when a side task is newly submitted to FreeRide, as shown in Figure 3, FreeRide’s automated side task profiler tracks its GPU memory consumption and per-step duration. During execution time, FreeRide employs one side task manager and multiple side task workers, one for each GPU. The side task manager assigns the newly submitted side task to one of the side task workers, which will create the side task process, based on the resulting profile. We instrument DeepSpeed to report the start timestamps and duration of bubbles to the side task manager that will initiate state transitions of each side task through remote procedure calls (RPCs) at the start and end of each bubble.
@@ 3.3. FreeRide Workflow
Putting the aforementioned ideas together, we present the workflow of FreeRide in Figure 3. First, programmers adapt their side task implementation using the interface provided by FreeRide (step ➊). FreeRide then automatically generates a profile of the side task’s characteristics (step ➋), which is submitted with the side task to the side task manager of FreeRide (step ➌). After the side task is submitted, based on the memory allocation of pipeline training and the characteristics of the side task, the side task manager will assign this side task to one of the workers (step ➍). When the main pipeline training workload is running, the side task manager continuously adds bubbles from the instrumented LLM training framework (step ➎); at the same time, it starts/pauses side tasks based on the available bubbles (step ➏).
Figure 4. State transitions in a side task program.
[…]
@@ 4.1. Programming Framework of FreeRide
Figure 4(a) describes the programming framework of a side task. The framework’s core is a state machine with five states and six state transitions. These five states capture the life cycle of a side task, from process creation to process termination, and correspond to different usage of hardware resources, e.g., GPU memory and GPU execution time. The six state transitions are used by the programmer to implement the user-defined logic of a side task. The programmer can override the state transition functions to customize their behavior, e.g., allocating or releasing hardware resources or performing computation on GPU. Once the side task is implemented, FreeRide automatically handles the state transitions at runtime. Next, we discuss the states and their transitions.
@@
$\bullet$ RUNNING. In this state, the side task executes the step-wise GPU workload. Referring to the example above of the model training side task, this step involves reading the next batch, computing the output and loss, updating the model weights, and resetting the optimizer states. The side task iteratively enters the RunNextStep() state transition to execute these steps until the side task manager transitions its state through PauseSideTask(). Therefore, in this state, the side task process uses both GPU memory and SMs.
$\bullet$ STOPPED. This state marks the end of the life cycle of a side task, where the side task process releases all of its hardware resources and terminates. It can be transitioned from states CREATED, PAUSED, and RUNNING through StopSideTask() initiated by the side task manager.
Figure 4(b) shows state transitions of a side task in Stage 0 of Figure 1. Initially, the side task is in the PAUSED (P) state. After four FP operations in the main training workload have finished, a bubble starts and the side task manager initiates StartSideTask() to transition the side task to the RUNNING (R) state. After the first bubble ends, the side task manager initiates PauseSideTask() to pause the side task. Then, the main training workload has BP operations and bubbles interleaved, leading to back-and-forth transitions between PAUSED and RUNNING states of the side task.
@@ 4.2. Interface for Side Task Implementation
Iterative programming interface. This is the preferred interface for side tasks in FreeRide. It periodically checks whether the side task manager has initiated any state transitions. If so, it executes the state transition functions in Figure 4(a) and updates the state of the side task. Then, if the side task is currently in the RUNNING state, it executes RunNextStep(). The programmer only has to override these transition functions to implement the side task. Pausing and resuming the side task, the transition of states, and communication with the FreeRide side task manager are all handled by the interface itself. GPU workloads that are naturally step-wise, e.g., model training, can be easily adapted to the iterative interface. We will discuss the adaption to this interface in Section 5 using an example.
Imperative programming interface. Not all side tasks can be explicitly implemented step-wise. Therefore, FreeRide provides the imperative interface as a fallback solution. The core is the function RunGpuWorkload() that allows the programmer to implement generic GPU side tasks without breaking them into steps. When the side task manager changes the state of the side task to RUNNING for the first time, the interface calls the RunGpuWorkload() function to execute the side task. The interface implements the pausing and resuming through signals (SIGTSTP and SIGCONT (job_control_signals, )) and calls StartSideTask() and PauseSideTask() inside the handlers of the two signals. The imperative interface offers better versatility at the cost of higher performance overhead (discussed in Section 5 and evaluated in Section 6.2).